User name Markowitz
Log entry time 00:01:00 on June13,2002
Entry number 85953
Started shift in controlled access during machine maintenance period. Mark Stevens and Scott Spiegel are in hall to fix read back of left arm Q1 helium flow readback sensor.
17:00 MCC calls to tell us they would like to steer beam to hall. Ask Mark and Scott to finish work and do access checklist.
18:00 Mark and Scott out of hall, MCC has brought us to beam permit.
18:30 Tuned up with 10 uA, checked that prescalers will be set at the higher current. Requested 110 uA.
Started run 1460, 677 K.
20:00 ROC14 died and had to be rebooted
Started run 1461, ~380K
20:15 TS1 died and had to be rebooted. While it was rebooting, Q3 tripped. It had a He level of 29% and would not reset. Phoned "tech-on-call" (Ed Folts) who is coming in to reset Q3, and requested an access from MCC. Richard Dawes in MCC related that HLA2 also crashed just before Q3. There is no current in Q3 and the other magnets are refilling but Q3 is still only at 30%.
Eventually Q3 got more helium and rebooted. We cycled Q3 to 1400 A then set it to the proper momentum. Apparently when the IOC rebooted the magnet JT was closed long enough trip off.
21:00 Started run 1462 239K, MCC took beam away at 21:27
21:45 started run 1463 700K
22:25 MCC takes beam to tune for Hall B
22:45 beam back.
24:00 Have about 300 deuterons at shift end